OpenStudy (anonymous):

Write the equation of the ellipse 9x2 + 4y2 – 72x + 40y + 208 = 0 in standard form.

OpenStudy (stamp):

\[9 (x-4)^2+4 (y+5)^2-36 = 0\]

OpenStudy (anonymous):

\[9(x-4)^2+4(y+5)^2-36=0\] \[9(x-4)^2+4(y+5)^2=36\] \[\frac{9(x-4)^2}{36} +\frac{4(y+5)^2}{36}=1\] \[\frac{(x-4)^2}{4}+\frac{(y+5)^2}{9}=1\]

OpenStudy (stamp):

yadig